Airphone, Inc., manufactures cellular telephones at a processing cost of $47 per unit. The company produces an average of 250 phones per week and has a yield of 87% good-quality phones, resulting in 13% defective phones, all of which can be reworked. The cost of reworking a defective telephone is $16.
a. Compute the quality-productivity ratio (QPR).

b. Compute the QPR if the company increased the production rate to 320 phones per week while reducing the processing cost to $42, reducing the rework cost to $12, and increasing the product yield of good-quality telephones to 94%.
Answer #1

1. FORMULA:

QPR = (GOOD UNITS / ((INPUT * PROCESSING COST) + (REWORKED UNITS * REWORK COST)) * 100

VALUES:

TOTAL PRODUCTION = 250
GOOD UNITS = 217.5
PROCESSING COST = 47
REWORKED UNITS = 32.5
REWORK COST = 16

QPR = (217.5 + 32.5) / ((250 * 47) + (32.5 * 16))) * 100= 2.04

2.

VALUES:

TOTAL PRODUCTION = 320
GOOD UNITS = 300.8
PROCESSING COST = 42
REWORKED UNITS = 19.2
REWORK COST = 16

QPR = (300.8 + 19.2) / ((320 * 42) + (19.2 * 16))) * 100= 2.33

**PLEASE RATE WITH A THUMBS UP :)